OpenStudy (anonymous):

one last one because i suck at absolute value. 2|3x-5|-8=8

OpenStudy (welshfella):

first add 8 to both sides then divide both sides by 2 you will be left with the equation |3x-5| = C where C is some number now split this up into 2 equations and solve them 3x - 5 = C and 3x - 5 = -C

OpenStudy (anonymous):

whats yo number?

OpenStudy (anonymous):

x = 13/3

OpenStudy (welshfella):

3x - 5 = 8 and 3x - 5 = -8

OpenStudy (welshfella):

x = 13/3 is one root

OpenStudy (welshfella):

what is the other?

OpenStudy (anonymous):

it has only one solution which is 13/3

OpenStudy (welshfella):

No the other solution is -1

OpenStudy (anonymous):

2(|3x−5|)−8=8 Add 8 to both sides. 2(|3x−5|)−8+8=8+8 2(|3x−5|)=16 Divide both sides by 2. 2(|3x−5|) / 2 = 16/2 |3x−5|=8 Solve Absolute Value. |3x−5|=8 We know either 3x−5=8 or 3x−5=−8 3x−5=8 3x−5+5=8+5 3x=13 3x/3 = 13/3 (Divide both sides by 3) x= 13 / 3 3x−5=−8 3x−5+5=−8+5(Add 5 to both sides) 3x=−3 3x/3 = −3 / 3 (Divide both sides by 3) x=−1 So ur answer is x = 13/3 or x = -1